# ACPO: AI-Enabled Compiler-Driven Program Optimization

### 15 Dec 2023 v1

@@ 19    UnrollingFactor = MLIF->getModelResultI("LU-Count");
A new ACPOModel class, namely ACPOLUModel can be instantiated as shown under Listing 1. We start by providing the necessary arguments for the constructor and initializing the private members. Subsequently, we run a few checks to make sure the steps mentioned prior under Section 3.6 are performed correctly. That leads us to the main inference method which is getAdviceML(). This method first retrieves a reference to the ML interface which is used to communicate with the ML framework. It then loads a model used for inference, which causes the framework to load the appropriate model as specified in the model.acpo file. If the model has already been loaded by the framework, then this command completes instantly, as loading the same model many times serves no purpose. The method then sets the inputs to the model and invokes the runModel method to begin inference.
The inference process takes place on the ML framework, and the runModel call is a blocking call that waits until the inference completes. When it does, a successful inference will return desired results, which can be retrieved using getModelResultI methods. However, if for any reason inference fails, the result of runModel will be false, indicating that the process failed. This mechanism allows the compiler to handle this case as the developer chooses. In our case, we forced the compiler to terminate with an error to ensure our ML model was the only mechanism for determining profitability. However, developers are welcome to provide fall-back mechanisms as they see fit.
To leverage ML in unrolling, we invoked the ACPOLUModel after the loop unrolling legality checks had been made. Also, we respected the high-priority overrides. The highest level is set by the explicit use of unroll-count by the user, followed by the usage of #pragma clang loop unroll_count(N). We invoke ACPOLUModel object subsequent to these two priorities and provide the unrolling information to the pass as shown in Listing 2.
@@ 4.2. ACPO Function Inlining
In this section, we show how to leverage ACPO to instantiate ACPOModel class for function inlining. Function inlining is one of the fundamental compiler optimizations used by many state-of-the-art compilers. In recent years, several attempts have been made to leverage ML on optimizing the pass for both code size reduction (Trofin et al., 2021; Theodoridis et al., 2022) and performance (Ashouri et al., 2022) and it was shown the optimal function inlining problem has an exponential increase in optimization space (Ashouri et al., 2018a; Theodoridis et al., 2022). A new ACPOModel class, namely, ACPOFIModel can be instantiated as shown under Listing 3 .
@@ 22                     ConstantInt::get(Type::getInt64Ty(*(getContextPtr())),
To leverage ML in function inlining, we invoked the ACPOFIModel by adding getACPOAdvice() after the legality checks have been made. Also, we respected the high-priority overrides. We invoke ACPOFIModel object and provide the inlining information to the pass as shown in Listing 4. We instantiate an ACPOFIModel instance, which in itself creates a persistent ML interface instance to a Python-based ML Framework, similar to what we showed earlier with loop unroll pass. Once the ACPOFIModel is instantiated, we invoke setCustomFeatures() to specify features as inputs to the model. This fulfills the requirements of the ACPOFIModel class and ensures all inputs are provided as needed before an inference is called. The getAdvice method is then called, which in turn invokes getAdviceML which facilitates the process of inference on behalf of the compiler. The result is then decomposed into ”FI-ShouldInline” field, which is then obeyed by the function inlining pass and applied as directed. Note that in this work, similar to loop unroll pass, we only replace the profitability checks with ML models.
@@ 5. Experimental Evaluation
In this Section, we provide details on our experimental campaign to showcase the efficacy of both ACPOLUModel by means of (1) standalone prediction accuracy, (2) cross-validation method using Polybench, and (3) the generalization evaluating of the model using a number of other standard benchmarks. Subsequently, we deploy and showcase the results we reproduced with ACPOFIModel.
5.1. ACPO Loop Unrolling Model
As mentioned earlier, Loop Unroll pass has been optimized from scratch by means of adding the necessary code to connect the interface with the ML side. Additionally, a new ML model has been designed and deployed for this matter. The details of the design and testing are provided as follows.
5.1.1. ACPOLUModel Architecture
Table 1 presents the architecture of ACPOLUModel. It is a 5-layer Neural Net with ReLu as the activation layer and a final Softmax layer to perform the classification task to identify the optimal unrolling factor. i.e., UP.Count values, given a loop. We have designed and implemented our model in PyTorch (Paszke et al., 2017), but once we train it, we convert the model into a TensorFlow (TF) (Abadi et al., 2016) frozen model and leverage TF’s inference mechanism to do inference at deployment. As mentioned earlier in Section 3.6, users can use any ML infrastructure of choice when training and deploying models as long as the input and output between the ML APIs and the compiler side match.
Table 1. ACPO Loop Unrolling Model Architecture
Layer No
Layer (type) Output Shape
0
Linear-FC1 [Features, 1, 32]
Relu-0 [Features, 1, 32]
1
Linear-FC1 [-1, 1, 128]
Relu-1 [-1, 1, 128]
2
Linear-FC2 [-1, 1, 256]
Relu-2 [-1, 1, 256]
3
Linear-FC3 [-1, 1, 64]
Relu-3 [-1, 1, 64]
4
Linear-FC4 [-1, 1, Classes]
Softmax [-1, 1, Classes]
The input to the model is a set of 30 handcrafted features, derived from the LLVM IR to characterize the state of any given loop being transformed. The features are passed with std::vector of pairs to the ML side. Additionally, the framework supports adding other types of features, i.e., graph-based, tokens, etc., however, in this work, we leverage handcrafted features derived from LLVM IR. The features are shown in Table 2. The Pearson (Benesty et al., 2009) correlation analysis between the features and the target label, i.e. classes corresponding to unrolling count, is depicted in Figure 3. The correlation values are within the range of [-1, +1], and specifically, we aimed at designing features with the highest positive/negative correlation with the target label. We did this by using features with at least a $\pm$5correlation, and this can be seen in the final column/row. We collect these features prior to the actual loop unrolling transformation routine and leverage setCustomFeatures() method to communicate with the ML side via the MLInterface API. Note that feature engineering and software characterization (Fursin et al., 2011; Ding et al., 2019; Cummins et al., 2021a) is a dynamic and fast-paced topic in the space and is outside the scope of this work. We leave the usage of other methods for our future work.
Table 2. ACPOLUModel Features
No Static Features No Static Features
1
PartialOptSizeThreshold
16
NumLoadInstPerLoopNest
2
AllowRemainder
17
NumStoreInstPerLoopNest
3
UnrollRemainder
18
TotLoopNestInstCount
4
AllowExpensiveTripCount
19
AvgNumLoadInstPerLoopNest
5
Force
20
AvgNumStoreInstPerLoopNest
6
TripCount
21
NumLoadInstPerLoop
7
MaxTripCount
22
NumStoreInstPerLoop
8
Size
23
TotLoopInstCount
9
InitialIVValueInt
24
AvgNumLoadInstPerLoop
10
FinalIVValueInt
25
AvgNumStoreInstPerLoop
11
StepValueInt
26
IsInnerMostLoop
12
NumPartitions
27
IsOuterMostLoop
13
IndVarSetSize
28
MaxLoopHeight
14
AvgStoreSetSize
29
TotBlocksPerLoop
15
AvgNumInsts
30
IsFixedTripCount
@@ 5.1.1. ACPOLUModel Architecture
5.1.2. Data Collection
We used the autotuner to collect a total of 10000 different unrolling configurations from Polybench by tuning the unroll count of individual loops (UP.Count), where the set of UP.Count values are $\in\{0,2,\cdots,64\}$. The autotuner uses the list of loop unrolling count opportunities to determine optimal configurations by using different search techniques. Then, the compiler uses the configurations suggested by the autotuner and generates a new executable file. We run the emitted executable and provide the execution times to the autotuner as feedback which will be leveraged by the autotuner to generate a new set of configurations based on the feedback. This process was repeated for 100 iterations to generate our training data. Note that each iteration, depending on the number of loops an application has, generates multiple data points, hence the total number of training data is significantly higher than the number of iterations of tuning rounds. The data generation process takes around 2 days.
The execution time of a number of Polybench applications can be too short to allow for a reliable measurement and to this end, we define a larger input dataset for those having an execution time of less than a second. This includes: gemver, jacobi-1d, bicg, atax, mvt, durbin, trisolv, gesummv, and deriche. In this work, we define the Loop Unrolling speedup as the following:
(1) $\footnotesize Loop_{speedup}=\frac{Loop_{runtime(Base)}}{Loop_{runtime(X)}}$
where $Loop_{runtime(Base)}$ is the execution time of a module having multiple loops at LLVM O3’s configuration and Loop_runtime(X), the same module having a loop unrolling count configuration of X in the set of all Unrolling configurations generated by our autotuner. In this work, total program runtime and the $T_{Loop}$ are collected with Linux perf. All the runtime measurements are done single-threaded, and we use numactl tool to bind the workloads to a unique CPU core. All experiments are measured on an ARMv8.2-A (Kunpeng 920) architecture @ 2.6GHz on Linux. We run each benchmark three times and before each measurement run, we flush the system page cache to avoid any perturbance into the collection of our training data and we repeat the process for all measurements having more than 1% of variance between the runs to make sure they are robust.
Additionally, we treat a single instance of iteration of data generation, normally having multiple loops and generating multiple samples, one set of runs meaning that we assign the Loop_speedup value to the whole set of rows generated as a result of the configuration. This way, we can filter out bad runs from good runs simply by pruning those causing a slowdown, or even having at least a certain amount of speedup values. This way ACPOLUModel only exposes to the set of quality samples and is trained to learn the relation between the given features and the target UP.Count.
5.1.3. Analysis of Compilation Time Overhead
@@
5.1.4. ML Training and Cross Validation
@@
The training is carried out using leave-one-out cross-validation for which we exclude one application from the training data and use it later as the test. This ensures no data leakage occurs. Table 4 showcases the classification performance of ACPOLUModel with its top-1 accuracy. As can be seen, ACPOLUModel is able to achieve a geometric mean of 80% accuracy with a standard deviation of 7% by means of its top-1 predictions. In certain applications, e.g., trmm, 2mm, etc., the top-1 accuracy is higher than the average, and in a few such as atax, it is lower. Our investigation reveals the reason lies in the imbalanced training set for the specific cases and it can be alleviated to some extent by employing a number of preprocessing and data augmentation techniques (Wei and Zou, 2019). We leave further investigation to our future works as they are outside the scope of this paper.
Table 4. ACPOLUModel Top-1 Accuracy on Polybench
Benchmark Top-1 Accuracy Benchmark Top-1 Accuracy
2mm 88.411 % gesummv 80 %
3mm 84.254 % gramschmidt 75.466 %
adi 76.543 % lu 85.971 %
atax 57.5 % ludcmp 85.938 %
bicg 63.295 % mvt 75.326 %
cholesky 86.414 % nussinov 79.734 %
correlation 84.091 % seidel-2d 80.909 %
covariance 86.42 % symm 79.797 %
deriche 82.481 % syr2k 84.858 %
doitgen 75.487 % syrk 82.84 %
floyd-warshall 82.864 % trisolv 69.225 %
gemver 86.547 % trmm 90.333 %
Geomean: 80.196%
Establishing a high standalone top-1 accuracy on unseen data is not the sole performance metric to evaluate a model and thus, we deployed the trained model using ACPO infrastructure on LLVM to measure the actual benefits of our ML model. Table 5 showcases the ACPOLUModel’s results when deployed on Polybench (Pouchet, 2012). We observe that on average, ACPOLUModel achieved a speedup of 4.07% among the Polybench benchmarks with a best of 49.3% on Floyd-Warshall. Additionally, we observe that out of 24 benchmarks, ACPOLUModel can outperform LLVM’s O3 on 15, excluding the close call for syr2k application.
Table 5. ACPOLUModel Speedup on Polybench
Benchmark   Time O3 Var O3 Size O3   Time ACPO Var ACPO Size ACPO Speedup Size Bloat
correlation   2.623 0.29 72728   2.739 0.984 72768 0.957 1.005
covariance   2.743 0.81 72648   2.734 0.927 72728 1.003 1.001
gramschmidt   3.98 0.94 72752   4.371 0.972 72752 0.910 1
cholesky   2.477 0.03 72752   2.248 0.027 72832 1.102 1.001
atax   0.739 0.26 72720   0.749 0.169 72696 0.985 0.999
2mm   3.478 0.97 72608   3.533 0.385 138216 0.984 1.903
3mm   5.182 0.078 72608   5.267 0.752 138248 0.983 1.9040
bicg   1.371 0.025 72664   1.343 0.85 72664 1.020 1
doitgen   0.897 0.044 72680   0.887 0.015 72688 1.011 1.0001
mvt   2.050 0.077 72608   1.959 0.085 72608 1.046 1
gesummv   1.499 0.003 72616   1.311 0.026 72616 1.143 1
gemver   2.699 0.15 72616   2.816 0.151 72648 0.958 1.0004
symm   2.728 0.33 72648   2.572 0.140 72680 1.060 1.0004
syr2k   3.150 0.91 72608   3.176 0.667 72680 0.991 1.0009
syrk   1.500 0.972 72608   1.496 0.879 72648 1.002 1.0005
trmm   1.378 0.92 72648   1.456 0.930 72680 0.946 1.0004
ludcmp   5.040 0.83 72720   4.812 0.6637 72784 1.047 1.0008
lu   5.186 0.31 72696   5.299 0.9422 72816 0.978 1.001
trisolv   1.146 0.013 72648   0.937 0.0171 72648 1.223 1
adi   18.70 0.017 72664   18.478 0.0012 72680 1.012 1.0002
seidel-2d   32.346 0.08 72648   32.318 0.005 72688 1.0008 1.0005
deriche   0.271 0.58 72648   0.251 0.978 72784 1.078 1.0018
floyd-Warshall   28.303 0.06 72624   18.946 0.0217 72656 1.493 1.0004
nussinov   3.839 0.08 72696   3.277 0.12 203808 1.171 2.803
Geomean   - 0.15 -   - 0.237 - 1.040725 1.10193
@@
Further investigation on our obtained results reveals that Floyd-Warshall application has a seven-loop module and the added loops from polybench.c with two instances of loop unrolling pass under O3 makes it for an opportunity set of 23 loop-nest regions which can be tuned by different unroll counts. ACPO’s prediction provides for a full unroll of a couple of loops including a multiple runtime unrolling of hot loops in the kernel which matches the optimal result our autotuner found after an exhaustive search. O3 on the other hand, only unrolled two loops by their available trip count of 3 and didn’t unroll the remaining hot loops. and thus, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops. Additionally, inspecting our training data generated by means of autotuning, we noticed there was an upper-bound of around 5-6% across Polybench by leveraging the optimal configuration for Loop Unrolling on Aarch64 and thus, achieving an average 4% (shown on Table 5), showcases the performance of the ACPOLUModel on the unseen application of Polybench in a cross-validation mode.
Table 6. ACPOLUModel Generalization Performance
Benchmarks   O3   ACPO   SpUp
Time Var Br Miss Size   Time Var Br Miss Size
[…]
5.1.5. ACPOLUModel Performance Generalization
In Section 5.1.4 we presented the cross-validation results on Polybench, where each application was used as a hold-out set (unseen test case) to evaluate the efficacy of the ACPOLUModel when trained with the remaining Polybench applications. Although the cross-validation results on Polybench show a promising trend for ACPOLUModel, we would still want our model to generalize well across a variety of completely unseen workloads. To this end, we choose a number of applications from Coral-2 benchmarks (Coral-2, 2017), Coremark (Gal-On and Levy, 2012), and Graph 500 (Murphy et al., 2010) to deploy ACPOLUModel on. We aim at selecting a diverse set of standard workloads covering a wide span of large-scale multi-grid solvers to the Embedded domain. Loop unrolling has a complex relation with other subsequent passes in the O3 pipeline, e.g., vectorization, etc, and thus, it is important for an ML model to be able to deduce the kind of correlation between the features derived by the compiler and the target performance metric. This time, we leverage all Polybench data in our training set and build an ACPOLUModel to deploy on the aforementioned benchmarks as test cases.
Table 6 demonstrates our result for which we gain a 5.4% speedup over CoreMark, an average 3% including an 18% speedup over AMG with Coral-2, and a 1.3% speedup on Graph500 SSSP. There are 3 instances, shown in bold under the Var column and by an *** in the Speedup column, that we were not able to contain the variance within 1% by means of multiple try-outs, and thus, we cannot reliably deduce a conclusion for them. Analyzing the obtained results, we notice that the majority of the performance comes from deciding the correct unrolling factor for the hot loops. For instance, Coral-2’s AMG runtime unrolling of a hot loop located at hypre_CSRMatrix class led to an increase in performance by almost 15% which O3 failed to unroll.
@@ 5.2.1. Function Inlining Model
MLGOPerf (Ashouri et al., 2022) is a reinforcement learning model trained with Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm. The model receives as input, 13 function features including block frequency, callsite height, caller and callee block counts and users, etc., and predicts whether or not a callsite is to be inlined. Readers can refer to the Table 2 of the reference paper for the exhaustive list of features used and the training methodology. The model has extended MLGO (Trofin et al., 2021)’s code size reduction to optimize performance instead. As shown in Listing 4, getACPOAdvice() initiates the call to collect necessary features and finally returns the Boolean value of the decision of whether or not to inline a callsite.

### 11 Mar 2024 v2

@@ 19    UnrollingFactor = MLIF->getModelResultI("LU-Count");
A new ACPOModel class, namely ACPOLUModel can be instantiated as shown under Listing 1. The textttgetAdviceML() method first retrieves a reference to the ML interface which is used to communicate with the ML framework. It then loads a model used for inference, which causes the framework to load the appropriate model as specified in the model.acpo file. The method then sets the inputs to the model and invokes the runModel method to begin inference. The inference process takes place on the ML framework, and the runModel call is a blocking call that waits until the inference completes. When it does, a successful inference will return desired results, which can be retrieved using getModelResultI methods. However, if for any reason inference fails, the result of runModel will be false, indicating that the process failed.
@@ 4.2. ACPO Function Inlining
In this section, we show how to leverage ACPO to instantiate ACPOModel class for function inlining. Function inlining is one of the fundamental compiler optimizations used by many state-of-the-art compilers. In recent years, several attempts have been made to leverage ML on optimizing the pass for both code size reduction (Trofin et al., 2021; Theodoridis et al., 2022) and performance (Ashouri et al., 2022) and it was shown the optimal function inlining problem has an exponential increase in optimization space (Ashouri et al., 2018; Theodoridis et al., 2022). A new ACPOModel class, namely, ACPOFIModel can be instantiated as shown under Listing 3 .
@@ 22                     ConstantInt::get(Type::getInt64Ty(*(getContextPtr())),
We invoked the ACPOFIModel by adding getACPOAdvice() after the legality checks have been made. Also, we respected the high-priority overrides. We invoke ACPOFIModel object and provide the inlining information to the pass as shown in Listing 4. We instantiate an ACPOFIModel instance, which in itself creates a persistent ML interface instance to a Python-based ML Framework, similar to what we showed earlier with loop unroll pass. Once the ACPOFIModel is instantiated, we invoke setCustomFeatures() to specify features as inputs to the model. This fulfills the requirements of the ACPOFIModel class and ensures all inputs are provided as needed before an inference is called. The getAdvice method is then called, which in turn invokes getAdviceML which facilitates the process of inference on behalf of the compiler. The result is then decomposed into ”FI-ShouldInline” field, which is then obeyed by the function inlining pass and applied as directed. Note that in this work, similar to loop unroll pass, we only replace the profitability checks with ML models.
@@ 5. Experimental Evaluation
In this Section, we provide details on our experimental campaign to showcase the efficacy of both ACPOLUModel by means of cross-validation method on Polybench Subsequently, we deploy and showcase the results we reproduced with ACPOFIModel.
ACPO Loop Unrolling Model
As mentioned earlier, Loop Unroll pass has been optimized from scratch by means of adding the necessary code to connect the interface with the ML side. Additionally, a new ML model has been designed and deployed for this matter. The details of the design and testing are provided as follows. ACPOLUModel is a 5-layer Neural Net with ReLu as the activation layer and a final Softmax layer to perform the classification task to identify the optimal unrolling factor. i.e., UP.Count values, given a loop. We have designed and implemented our model in PyTorch (Paszke et al., 2017), but once we train it, we convert the model into a TensorFlow (TF) (Abadi et al., 2016) frozen model and leverage TF’s inference mechanism to do inference at deployment. As mentioned earlier in Section 3.5, users can use any ML infrastructure of choice when training and deploying models as long as the input and output between the ML APIs and the compiler side match. The input to the model is a set of 30 handcrafted features, derived from the LLVM IR to characterize the state of any given loop being transformed. The features are passed with std::vector of pairs to the ML side. Additionally, the framework supports adding other types of features, i.e., graph-based, tokens, etc., however, in this work, we leverage handcrafted features derived from LLVM IR. The features used are shown in Table 3 and taken from our super class of ACPOFeatures which contains a large set of handcrafted features defined in different scopes, i.e., Module, Function, Call Graph, and Loops, so that users can easily add and register them into their optimization use cases. The Pearson (Benesty et al., 2009) correlation analysis between the features and the target label, i.e. classes corresponding to unrolling count, is also depicted in Figure 3. The correlation values are within the range of [-1, +1], and specifically, we aimed at designing features with the highest positive/negative correlation with the target label. We did this by using features with at least a $\pm$2correlation.
@@
Data Collection
We used the autotuner to collect a total of 10000 different unrolling configurations from Polybench by tuning the unroll count of individual loops (UP.Count), where the set of UP.Count values are $\in\{0,2,\cdots,64\}$. The autotuner uses the list of loop unrolling count opportunities to determine optimal configurations by using different search techniques. Then, the compiler uses the configurations suggested by the autotuner and generates a new executable file. This process was repeated for 100 iterations to generate our training data. Note that each iteration, depending on the number of loops an application has, generates multiple data points, hence the total number of training data is significantly higher than the number of iterations of tuning rounds. The data generation process takes around 2 days. The execution time of a number of Polybench applications can be too short to allow for a reliable measurement and to this end, we define a larger input dataset for those having an execution time of less than a second. This includes: gemver, jacobi-1d, bicg, atax, mvt, durbin, trisolv, gesummv, and deriche. In this work, we define the Loop Unrolling speedup as the following:
(1) $\scriptsize Loop_{speedup}=\frac{Loop_{runtime(Base)}}{Loop_{runtime(X)}}$
where $Loop_{runtime(Base)}$ is the execution time of a module having multiple loops at LLVM O3’s configuration and Loop_runtime(X), the same module having a loop unrolling count configuration of X in the set of all Unrolling configurations generated by our autotuner. In this work, total program runtime and the $T_{Loop}$ are collected with Linux perf. All the runtime measurements are done single-threaded, and we use numactl tool to bind the workloads to a unique CPU core. All experiments are measured on an ARMv8.2-A (Kunpeng 920) architecture @ 2.6GHz on Linux. We run each benchmark three times and before each measurement run, we flush the system page cache to avoid any perturbance into the collection of our training data and we repeat the process for all measurements having more than 1% of variance between the runs to make sure they are robust. Additionally, we treat a single instance of iteration of data generation, normally having multiple loops and generating multiple samples, one set of runs meaning that we assign the Loop_speedup value to the whole set of rows generated as a result of the configuration. This way, we can filter out bad runs from good runs simply by pruning those causing a slowdown, or even having at least a certain amount of speedup values. This way ACPOLUModel only exposes to the set of quality samples and is trained to learn the relation between the given features and the target UP.Count.
Analysis of Compilation Time Overhead
@@
ML Training and Cross Validation
@@
The training is carried out using leave-one-out cross-validation for which we exclude one application from the training data and use it later as the test. This ensures no data leakage occurs. Measured Top-1 accuracy of the ACPOLUModel’s predictions exhibits a geometric mean of 80% accuracy with a standard deviation of 7% on Polybench. In certain applications, e.g., trmm, 2mm, etc., the top-1 accuracy is higher than the average, and in a few such as atax, it is lower. Our investigation reveals the reason lies in the imbalanced training set for the specific cases and it can be alleviated to some extent by employing a number of preprocessing and data augmentation techniques (Wei and Zou, 2019). We leave further investigation to our future works as they are outside the scope of this paper. Subsequent to establishing the high top-1 accuracy on a cross validation mode, we evaluate the performance of the ACPOLUModel by means of deploying it on Polybench and measuring the actual benefits of our ML model. Table 2 showcases the ACPOLUModel’s results when deployed on Polybench (Pouchet, 2012) on Aarch64. We observe that on average, ACPOLUModel achieved a speedup of 4.07% among the Polybench benchmarks with a best of 49.3% on Floyd-Warshall. Additionally, we observe that out of 24 benchmarks, ACPOLUModel can outperform LLVM’s O3 on 15, excluding the close call for syr2k application.
@@
Further investigation on our obtained results reveals that Floyd-Warshall application has a seven-loop module and the added loops from polybench.c with two instances of loop unrolling pass under O3 makes it for an opportunity set of 23 loop-nest regions which can be tuned by different unroll counts. ACPO’s prediction provides for a full unroll of a couple of loops including a multiple runtime unrolling of hot loops in the kernel which matches the optimal result our autotuner found after an exhaustive search. O3 on the other hand, only unrolled two loops by their available trip count of 3 and didn’t unroll the remaining hot loops. and thus, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops.
@@ 5.1.1. Function Inlining Model
MLGOPerf (Ashouri et al., 2022) is a reinforcement learning model trained with Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm. The model receives as input, 13 function features including block frequency, callsite height, block counts, etc., and predicts whether or not a callsite is to be inlined. Readers can refer to the Table 2 of the reference paper for the exhaustive list of features used and the training methodology. The model has extended MLGO (Trofin et al., 2021)’s code size reduction to optimize performance instead. As shown in Listing 4, getACPOAdvice() initiates the call to collect necessary features and finally returns the Boolean value of the decision of whether or not to inline a callsite.